OpenStudy (anonymous):

There are an unlimited amount of toy cars that come in the colors red, blue, yellow and green. If a kid receives three cars at random, what is the probability that two will be the same color and one will be different?

OpenStudy (anonymous):

kind of weird since we don't know the ratios of the numbers of cars just because there are an infinite amount of each, doesn’t mean they are in equal proportion does it?

OpenStudy (anonymous):

we can pretend they are i guess, but it doesn't say it

OpenStudy (anonymous):

how might you go about setting up an equation for something like this? i think i'm getting thrown off by the fact that they are of unlimited quantity and i don't have any numbers to begin from.

OpenStudy (anonymous):

that is the problem you can't really do it without assuming they are in equal proportion, one third of each color if you get to assume that, then i guess we can do it without too much difficulty (famous last words)

OpenStudy (anonymous):

if we assume they are in equal proportion, then the probability of any one particular pick is \[\frac{1}{3}\times \frac{1}{3}\times \frac{1}{3}=\frac{1}{27}\] then we can count how many ways to pick 2 of one and one of the other it is not that many (only 6) so we can list them

OpenStudy (anonymous):

for example RR B BB R RR Y YY R BB Y YY B

OpenStudy (anonymous):

unless i missed any

OpenStudy (anonymous):

oh damn, that is wrong, sorry order doesn't count lets to it an easier way

OpenStudy (anonymous):

the probability they are all blue is \(\frac{1}{27}\) same for all yellow and all red we can now compute \[1-\frac{3}{27}\]

OpenStudy (anonymous):

why do we say 1/27 as the probability instead of 1/3

OpenStudy (anonymous):

for example all blue first one is blue and second one is blue and third one is blue \[\frac{1}{3}\times \frac{1}{3}\times \frac{1}{3}\]

OpenStudy (anonymous):

same for all red and all yellow

OpenStudy (anonymous):

therefore probability of all blue is \(\frac{1}{27}\) probability of all red is \(\frac{1}{27}\) probability of all yellow is \(\frac{1}{27}\) since these are clearly disjoint events, the probability that they are any one color is \[\frac{1}{27}+\frac{1}{27}+\frac{1}{27}=\frac{3}{27}=\frac{1}{9}\]

OpenStudy (anonymous):

and so the probability you get two different colors is \[1-\frac{1}{9}\] if my reasoning is correct

OpenStudy (anonymous):

do you mean two of the same color?

OpenStudy (anonymous):

oh this makes sense. nevermind! thank you very much!

OpenStudy (anonymous):

should i multiply by on more though since there are four colors?